OpenStudy (anonymous):

What is the solution of the system? 2x – y = –7 4x – y = –4

OpenStudy (anonymous):

So, can I teach you how to use wolfram alpha if you just want answers? It's faster and more accurate than asking people here.

OpenStudy (anonymous):

1) Go to http://www.wolframalpha.com

OpenStudy (anonymous):

2) Type this: 2x-7 = -7, 4x - y = -4

OpenStudy (anonymous):

3) Press return

OpenStudy (anonymous):

oh yeah cuz i went on there and they gave me the wrong answer but thank you very much i appreciate it

OpenStudy (anonymous):

What did you do? That's how I'm giving you the answers.

OpenStudy (anonymous):

You have to put a comma between the two equations

OpenStudy (anonymous):

oh noes wolfram gives incorrect answers!!!

OpenStudy (anonymous):

2x – y = –7 (1) 4x – y = –4 (2) negate the second equation -4x + y = 4 (3) Add (1) and (3) -2x + 0 = -3 => x = 3/2 Substitute this in (1) to get y 3 - y = -7 => -y = -10 => y = 10